Marx Healthcare, a group practice clinic with 10 physicians, had the following income in 2017:

Revenue $3,350,000
Less operating expenses:
Salaries
Physicians $1,340,000
Nurses 146,000
Nursing aide 62,500
Receptionist 50,000
Accounting services 46,000
Training 130,000
Supplies 259,000
Phone and fax 3,850
Insurance 260,000
Depreciation 230,000
Utilities 18,600
Miscellaneous 68,000
Total operating expenses 2,613,950
Income before taxes 736,050
Less taxes on income 257,618
Net income $478,432


The following changes are expected in 2018:

1. The clinic is expecting a 2 percent decline in revenues because of increasing pressure from insurance companies.
2. Physicians are planning to hire a physician assistant at a salary of $46,000 per year.
3. Training costs are expected to increase by $10,000.
4. Supplies are expected to increase to be 10 percent of revenue.
5. Phone, fax, accounting services and insurance amounts will stay the same.
6. Depreciation expense will increase by $22,000 per year, since the clinic is planning to purchase equipment for $125,000.
7. Utilities and miscellaneous expenses are expected to increase by 5 percent next year.
8. Taxes on income will be 35 percent.


Prepare a budgeted income statement for Marx Healthcare for the year 2018. (Round answers to 0 decimal places, e.g. 125.)
Answer #1

Effect of changes:

1)As the revenue decrease by 2% then revenue becomes

3350000-2%=3283000

2)Hence there will be increase in salary expense.

3)training costs increase by 10000 then it becomes 130000+10000=140000

4)supplies expect to increase by 10% of revenue.hence

3283000*10%=328300

5)phone fax accounting services remain same hence no change in them.

6)depreciation expense increase by 22000 hence depreciation becomes 252000

7)utilities and miscellaneous expense increase by 5% hence

86600+5%=90930.

8)taxes on income shall be 35%.

Hence income statement shall be

Revenue:3283000

Salaries consolidated:(1644500)

Training costs:(140000)

Supplies:(328300)

Phone,fax&accounting services: (39850)

Depreciation:(252000)

Utilities and misc.exp:(90930)

Income before taxes:787420

(-)taxes@35% :(275597)

Net income . :511823